# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ III-A2 Max strategy
A local center $c_{i}$ is considered isolated if, for all $c_{j}\neq c_{i}$, the similarity between the neighborhoods associated with $c_{i}$ and $c_{j}$ is zero, i.e.,
$\mathbb{M}_{ij}=0.$
In other words, an isolated local center is one whose neighborhood does not overlap with any other neighborhoods. As the max strategy assumes each cluster consists of various convex shapes, if there are any isolated local centers, they will be discarded.
@@
An everywhere non-convex cluster is characterized by all of its points having non-convex neighborhoods. Examples include one-dimensional manifolds, such as spirals, and self-crossing curves. A region containing such a cluster is termed a locally non-convex region.
The max strategy fails to identify everywhere non-convex clusters because as the depth functions used herein only suit convex shapes, preventing recognition of local centers satisfying $r_{i}(\mathbf{x}_{i})\leq 2$ in a non-convex neighborhood. To assess the presence of everywhere non-convex clusters and mitigate their impact, we categorize local centers into three sets based on the outcome of prior filtering:
  • Set $\mathbf{c}_{1}$ includes all local centers that remain after initial filtering, characterized by non-isolated local centers with $r_{i}(\mathbf{x}_{i})\leq 2$ and $f\neq 1$.
  • Set $\mathbf{c}_{2}$ contains non-isolated local centers $c_{v}$ for which $\forall c_{i}\in\mathbf{c}_{1}$, $c_{v}$ satisfies $\mathbb{M}_{vi}=0$.
  • Set $\mathbf{c}_{3}$ consists of local centers $c_{v}\notin\mathbf{c}_{1}$ satisfying $\exists c_{i}\in\mathbf{c}_{1}$ such that $\mathbb{M}_{vi}>0$.
The presence of everywhere non-convex clusters is inferred if $\mathbf{c}_{2}$ is not empty. In such cases, Algorithm 4 are taken to segregate locally non-convex regions from other clusters.
@@
1:Sets of local centers $\mathbf{c}_{1}$, $\mathbf{c}_{2}$, and $\mathbf{c}_{3}$, neighborhood similarity matrix $\mathbb{M}$
@@
3:if $\mathbf{c}_{2}\neq\emptyset$ then
4:     for $v$ in the indices of local centers in $\mathbf{c}_{3}$ do
5:         if  $\max_{j:c_{j}\in\mathbf{c}_{2}}\mathbb{M}_{vj}>\max_{i:c_{i}\in\mathbf{c}_{1}}\mathbb{M}_{vi}$  then
6:              $\mathbf{c}_{2}=\mathbf{c}_{2}\cup\{c_{v}\}$
7:              for $i$ in the indices of local centers in $\mathbf{c}_{1}$ do
8:                  if $\mathbb{M}_{vi}>0$ then
9:                       $\mathbf{c}_{1}=\mathbf{c}_{1}\setminus\{c_{i}\}$
@@
14:     $\mathbf{C}=\mathbf{c}_{1}\cup\mathbf{c}_{2}$
15:     if any isolated local centers exist based on $\mathbf{C}$ then
@@
Algorithm 4 ensures that local centers in $\mathbf{c}_{2}$, which capture the essence of locally non-convex regions, remain disjoint with the local centers in other clusters, i.e., local centers in $\mathbf{c}_{1}$.
@@ III-B Clustering in DLCC
Assuming the grouping results for the filtered centers are at hand, with the unique neighbors of each group serving to initialize temporary clusters, let $\mathbf{C}$ represent the set of filtered centers, such that $\mathbf{C}=\bigcup_{k=1}^{G}\mathbf{g}_{k}$, where $G$ denotes the total number of groups, and $\mathbf{g}_{k}$ contains the filtered centers of the $k$th group. Consequently, the number of clusters $K$ is inherently defined to be equal to $G$, as each group’s unique neighbors directly initiate a corresponding temporary cluster. The subsequent step involves defining a score for each observation with respect to each cluster in order to update these temporary clusters.
III-B1 Min strategy
@@
For the min strategy, the score accounts for both clustering and assessing if each observation’s clustering result is acceptable, and it is computed using the similarity matrix $\mathbb{S}$. If $\mathbb{S}$ is asymmetric, it is updated by replacing $\mathbb{S}_{ji}$ and $\mathbb{S}_{ij}$ with their average value. For clarity, we continue to denote the symmetrized matrix as $\mathbb{S}$, ensuring that $\mathbb{S}_{ij}=\mathbb{S}_{ji}$ in this and the subsequent discussion about the DC metric in Section IV. We define the score function for an observation $\mathbf{x}_{i}$ with respect to a temporary cluster $k$ under the min strategy as
$\operatorname{score}_{i|k}=\frac{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji}-\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}}{\max\{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji},\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}\}\,},$ (16)
@@
where $\mathcal{I}_{k}$ is the set of indices for filtered centers within the $k$th group, i.e., $\mathcal{I}_{k}=\{j:\mathbf{x}_{j}\in\mathbf{g}_{k}\}$. The score (16) takes a value between $-1$ and $1$ and a positive value indicates that, compared with the nearest filtered center in other groups, the similarity between $\mathbf{x}_{i}$ and the nearest filtered center in group $k$ is larger.
For each cluster, two score pools are generated: $\mathcal{S}_{k}$, containing both the unique neighbors within the cluster ($\mathbf{U}_{\mathbf{g}_{k}}$) and their corresponding scores, and $\mathcal{\hat{S}}_{k}$, holding observations not in $\mathbf{U}_{\mathbf{g}_{k}}$ yet having positive scores toward this cluster, along with their scores. Upon reviewing these pools, observations with high scores may shift from $\mathcal{\hat{S}}_{k}$ to $\mathcal{S}_{k}$, and vice versa, for those with low scores in $\mathcal{S}_{k}$. Lastly, $\mathcal{S}_{k}$ includes observations confidently assigned to the cluster, while those in $\mathcal{\hat{S}}_{k}$ remain unlabelled. Note that during the reassignment of observations between pools, their corresponding scores are also moved. The detailed steps are outlined as Algorithm 5, where the second element of the min function in line $14$ ensures at least $s/2$ observations per temporary cluster, preventing overly small clusters. Figure 4 illustrates an example using the Yale B face recognition dataset [56]. Algorithm 5 removes labels from unique neighbors within each group that are less distinguishable from other points in their own group. Meanwhile, it assigns labels to points that are not unique neighbors but are relatively closer to the filtered centers of their group.
Figure 4: t-SNE visualization [47] of the Yale B dataset. Colors represent the true labels of the points, with only those in temporary clusters being colored. Points enclosed by diamond-shaped boxes represent filtered centers. (a) Initial points in $\mathcal{S}_{k}$, with points enclosed by square boxes indicating those to be moved to $\mathcal{\hat{S}}_{k}$. (b) The temporary clusters obtained after applying Algorithm 5.
Algorithm 5 Update Temporary Cluster
1:Score pools $\mathcal{S}_{k}$ and $\hat{\mathcal{S}}_{k}$ for $k=1,\ldots,K$, where $\mathcal{S}_{k}^{1}$ and $\hat{\mathcal{S}}_{k}^{1}$ represent observations, and $\mathcal{S}_{k}^{2}$ and $\hat{\mathcal{S}}_{k}^{2}$ represent the corresponding scores, neighborhood size $s$.
2:Updated score pool $\mathcal{S}_{k}$, where $\mathcal{S}_{k}^{1}$ constructs the $k$-th temporary cluster.
3:for $k=1$ to $K$ do
4:     $\gamma_{1}=\text{mean}(\hat{\mathcal{S}}_{k}^{2})$
5:     for $i$ in $\mathcal{S}_{k}^{2}$ do
6:         if $i<\gamma_{1}$ then
7:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\setminus\{i\}$
8:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\cup\{i\}$
9:              Move corresponding points from $\mathcal{S}_{k}^{1}$ to $\hat{\mathcal{S}}_{k}^{1}$
10:         end if
11:     end for
12:     $SS=\text{Sort scores in }\mathcal{S}_{k}^{2}\text{ in descending order}$
13:     $B=\text{length of }SS$
14:     $idx=\arg\max_{j=\text{floor}(B/2)\text{ to }B-1}(SS[j]-SS[j+1])$
15:     $\hat{B}=\text{length of }\hat{\mathcal{S}}_{k}^{2}$
16:     $\gamma_{2}=\min(SS[idx],1-\frac{s/2-B}{\hat{B}}\text{quantile of }\hat{\mathcal{S}}_{k}^{2})$
17:     for $u$ in $\hat{\mathcal{S}}_{k}^{2}$ do
18:         if $u>\gamma_{2}$ then
19:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\setminus\{u\}$
20:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\cup\{u\}$
21:              Move corresponding points from $\hat{\mathcal{S}}_{k}^{1}$ to $\mathcal{S}_{k}^{1}$
22:         end if
23:     end for
24:end for
III-B2 Max strategy
The score for the max strategy is only for overlapping neighbors between filtered centers in different groups, and is defined as
$\operatorname{score}_{i|k}=\frac{|\{c\in\mathbf{g}_{k}\mid\mathbf{x}_{i}\in\mathbf{N}_{c}\}|}{|\mathbf{g}_{k}|}.$ (17)
Observations that are neighbors of filtered centers in different groups are allocated to the cluster with the highest score. Unlabeled observations either do not have the highest unique score or are not neighbours of any filtered center.
III-B3 Classification
Among three built-in classification methods for the remaining unlabelled observations, MDC is the non-parametric method based on data depth. MDC assigns each observation to the cluster within which it achieves its maximum depth value. To ensure consistency in our methodology, we employ the same type of data depth in the MDC as we used in defining the local centers. When Mahalanobis depth is chosen, the covariance matrix in the classification step is estimated using traditional moment estimation methods. This is based on the theoretical consideration that temporary clusters are subsets of clusters and all observations in these temporary clusters are selected with a degree of confidence.
As for the KNN method, it is an intuitive classification method and is suitable for non-convex shapes. Given that we have already constructed the similarity matrix $\mathbb{S}$, the KNN method in our algorithm identifies neighbors through depth-based similarity, rather than relying on Euclidean distance. As such, the method identifies neighbors directly within $\mathbb{S}$, thus eliminating the need for additional computational steps.
The last method, RF, is a renowned machine learning algorithm and has a wide range of applications. In classification, it operates without assumptions on the shapes of the clusters and can demonstrate satisfactory performance even in high-dimensional spaces.
IV Exploring Parameter Selections

### v2

@@ III-A2 Max Strategy
A local center $c_{i}$ is considered isolated if, for every other local center $c_{j}\neq c_{i}$, the neighborhood similarity satisfies
$\mathbb{M}_{i,j}=0.$
Here, $\mathbb{M}$ denotes the neighborhood similarity matrix over local-center candidates, where $\mathbb{M}_{i,j}$ records the similarity between the neighborhoods associated with $c_{i}$ and $c_{j}$.
In other words, an isolated local center is one whose neighborhood does not overlap with any other neighborhoods. As the max strategy assumes each cluster consists of various convex shapes, isolated local centers are discarded to avoid defining extremely small clusters that may in fact correspond to outliers.
@@
An everywhere non-convex cluster is characterized by all of its points having non-convex neighborhoods.
Examples of everywhere non-convex cluster include one-dimensional manifolds, such as spirals, and self-crossing curves. A region containing such a cluster is termed a locally non-convex region. The max strategy fails to identify everywhere non-convex clusters because as the depth functions used herein only suit convex shapes, preventing recognition of local centers satisfying $r_{i}(\mathbf{x}_{i})\leq 2$ in a non-convex neighborhood. To assess the presence of everywhere non-convex clusters and mitigate their impact, we categorize local centers into three sets based on the outcome of prior filtering:
  • Set $\mathcal{C}_{1}$ includes all local centers that remain after initial filtering, characterized by non-isolated local centers with $r_{i}(\mathbf{x}_{i})\leq 2$ and $f\neq 1$.
  • Set $\mathcal{C}_{2}$ contains non-isolated local centers $c_{v}$ for which $\forall c_{i}\in\mathcal{C}_{1}$, $c_{v}$ satisfies $\mathbb{M}_{v,i}=0$.
  • Set $\mathcal{C}_{3}$ consists of local centers $c_{v}\notin\mathcal{C}_{1}$ satisfying $\exists c_{i}\in\mathcal{C}_{1}$ such that $\mathbb{M}_{v,i}>0$.
The presence of everywhere non-convex clusters is inferred if $\mathcal{C}_{2}$ is not empty. In such cases, Algorithm 4 is used to segregate locally non-convex regions from other clusters.
@@
1:Sets of local centers $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, and $\mathcal{C}_{3}$, neighborhood similarity matrix $\mathbb{M}$
@@
3:if $\mathcal{C}_{2}\neq\emptyset$ then
4:  for $v$ in the indices of local centers in $\mathcal{C}_{3}$ do
5:    if $\max_{j:c_{j}\in\mathcal{C}_{2}}\mathbb{M}_{v,j}>\max_{i:c_{i}\in\mathcal{C}_{1}}\mathbb{M}_{vi}$ then
6:     $\mathcal{C}_{2}=\mathcal{C}_{2}\cup\{c_{v}\}$
7:     for $i$ in the indices of local centers in $\mathcal{C}_{1}$ do
8:      if $\mathbb{M}_{v,i}>0$ then
9:         $\mathcal{C}_{1}=\mathcal{C}_{1}\setminus\{c_{i}\}$
@@
14:  $\mathbf{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$
15:  if any isolated local centers exist within the updated set $\mathbf{C}$ then
@@
This Algorithm ensures that local centers in $\mathcal{C}_{2}$, which capture the essence of locally non-convex regions, remain disjoint with the local centers in other clusters, i.e., local centers in $\mathcal{C}_{1}$.
@@ III-B Clustering in DLCC
The grouping results $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{K}\}$ for the filtered centers are at hand and $\mathbf{C}$ represent the set of filtered centers. Consequently, the number of clusters is inherently defined to be equal to $K$, as each group’s unique neighbors directly initiate a corresponding temporary cluster. The subsequent step involves defining a score for each observation w.r.t. each cluster in order to update these temporary clusters.
@@
The score accounts for both clustering and assessing if each observation’s clustering result is acceptable, and it is computed using the similarity matrix $\mathbb{S}$. Since $\mathbb{S}$ is asymmetric, it is updated by replacing $\mathbb{S}_{j,i}$ and $\mathbb{S}_{i,j}$ with their average value. For clarity, we continue to denote the symmetrized matrix as $\mathbb{S}$, ensuring that $\mathbb{S}_{i,j}=\mathbb{S}_{j,i}$. We define the score function for an observation $\mathbf{x}_{i}$ w.r.t. a temporary cluster $k$ under as
$\operatorname{score}_{i|k}=\frac{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{j,i}-\max_{j\in\bigcup_{z\neq k}\mathcal{I_{z}}}\mathbb{S}_{j,i}}{\max\{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{j,i},\max_{j\in\bigcup_{z\neq k}\mathcal{I_{z}}}\mathbb{S}_{j,i}\}\,}$, (16)
@@
where $\mathcal{I}_{k}$ is the set of indices for filtered centers within the $k$th group, i.e., $\mathcal{I}_{k}=\{j:\mathbf{x}_{j}\in\mathbf{g}_{k}\}$. The score (16) takes a value between 1-1 and $1$ and a positive value indicates that, compared with the nearest filtered center in other groups, the similarity between $\mathbf{x}_{i}$ and the nearest filtered center in group $k$ is larger.
For each cluster, two pairs of sets are generated: $\mathbf{T}_{k}$, containing the unique neighbors within the cluster ($\mathbf{U}_{\mathbf{g}_{k}}$), and $\mathcal{S}_{k}$, the corresponding scores for points in $\mathbf{T}_{k}$ w.r.t. the $k$th cluster; as well as $\hat{\mathbf{T}}_{k}$, holding unassigned observations with positive scores toward the $k$th cluster, and $\hat{\mathcal{S}}_{k}$, their associated scores. During the update process, observations in $\hat{\mathbf{T}}_{k}$ with sufficiently high scores in $\hat{\mathcal{S}}_{k}$ may be moved to $\mathbf{T}_{k}$ and $\mathcal{S}_{k}$, while those in $\mathbf{T}_{k}$ with low scores in $\mathcal{S}_{k}$ may be moved to $\hat{\mathbf{T}}_{k}$ and $\hat{\mathcal{S}}_{k}$. After refinement, $\mathbf{T}_{k}$ contains the observations confidently assigned to the $k$th cluster, while $\hat{\mathbf{T}}_{k}$ collects those remaining unassigned. We refer to $\{\mathbf{T}_{k}\}_{k=1}^{K}$ as the set of temporary clusters.
The detailed steps are outlined in Algorithm 5, with only a minor difference between the min and max strategies. Under the max strategy, we use a single mean-based cutoff for each cluster, applied uniformly to points in $\mathbf{T}_{k}$ and $\hat{\mathbf{T}}_{k}$ and take $\gamma=0.5$, yielding a looser threshold for admitting points into temporary clusters. Under the min strategy, the additional quantile-based cap in Line 9 enforces at least $s/2$ observations per temporary cluster, preventing overly small clusters.
Figure 4 presents an example based on the Yale B face recognition dataset [44]. Algorithm 5 updates the temporary clusters by discarding unique neighbors whose scores relative to their currently assigned clusters are insufficient, while simultaneously assigning labels to non–unique neighbors that are sufficiently close to the filtered centers of their respective groups. As shown in the figure, this procedure removes visually misclassified points while correctly labeling most points near local centers.
Figure 4: t-SNE visualization [45] of the Yale B dataset in two-dimensional space (see Section V for dataset description). Colors indicate the temporary labels of the points, i.e., membership in $\{\mathbf{T}_{k}\}_{k=1}^{K}$. From left to right, the panels illustrate the progression from the initial state (unique neighbors) to the temporary clusters after Algorithm 5. Points enclosed by diamond-shaped boxes denote filtered centers.
Algorithm 5 Updating Temporary Clusters
1:For $k=1,\ldots,K$, temporary clusters $\mathbf{T}_{k}$, scores $\mathcal{S}_{k}$ for points in $\mathbf{T}_{k}$; Remaining observations for each cluster $\hat{\mathbf{T}}_{k}$ and their scores $\hat{\mathcal{S}}_{k}$; neighborhood size $s$; $\texttt{strategy}\in\{\texttt{min},\texttt{max}\}$.
2:Updated $\mathbf{T}_{k}$ and $\mathcal{S}_{k}$ for $k=1,\ldots,K$.
3:Set $\gamma\leftarrow 1$ if $\texttt{strategy}=\texttt{min}$, otherwise $\gamma\leftarrow 0.5$.
4:for $k=1$ to $K$ do
5:  $\alpha_{1}\leftarrow\gamma\cdot\mathrm{mean}(\hat{\mathcal{S}}_{k})$ $\triangleright$ if $\hat{\mathcal{S}}_{k}=\emptyset$, use $\mathrm{mean}(\hat{\mathcal{S}}_{\mathrm{all}})$, where $\hat{\mathcal{S}}_{\mathrm{all}}=\bigcup_{\ell:\hat{\mathcal{S}}_{\ell}\neq\emptyset}\hat{\mathcal{S}}_{\ell}$
6:  for each $i\in\mathbf{T}_{k}$ with score $\alpha_{1}$ do
7:    $\mathbf{T}_{k}\leftarrow\mathbf{T}_{k}\setminus\{i\}$,   $\hat{\mathbf{T}}_{k}\leftarrow\hat{\mathbf{T}}_{k}\cup\{i\}$; update $\mathcal{S}_{k}$ and $\hat{\mathcal{S}}_{k}$.
8:  end for
9:  if strategy = min then
10:    $SS\leftarrow$ sorted scores in $\mathcal{S}_{k}$ (descending)
11:    $idx\leftarrow\arg\max_{j=\mathrm{floor}(|\mathcal{S}_{k}|/2)\text{ to }|\mathcal{S}_{k}|-1}(SS[j]-SS[j+1])$
12:    $\alpha_{2}\leftarrow\min(SS[idx],1-\frac{s/2-|\mathcal{S}_{k}|}{|\hat{\mathcal{S}}_{k}|}$th quantile of $\hat{\mathcal{S}}_{k})$
[…]
14:    $\alpha_{2}\leftarrow\alpha_{1}$.
15:  end if
16:  for each $u\in\hat{\mathbf{T}}_{k}$ with score $\hat{\alpha}_{2}$ do
17:    $\hat{\mathbf{T}}_{k}\leftarrow\hat{\mathbf{T}}_{k}\setminus\{u\}$,   $\mathbf{T}_{k}\leftarrow\mathbf{T}_{k}\cup\{u\}$.
18:  end for
19:end for
III-B1 Classification
After obtaining temporary clusters, we further assign the remaining unlabeled observations via a classification step. The motivation is that, for exemplar-based clustering methods, assigning ambiguous points solely according to their similarity to an exemplar can be unreliable, as such points are often weakly associated with any single exemplar. Instead, incorporating information from confidently assigned samples provides a more robust basis for assigning these ambiguous points. This idea has also been explored in other clustering researches. For example, Shi et al. [46] noted that, beyond center-based allocation, assignments of determined samples can be used to infer labels of undetermined ones.
Two built-in options for classifying the remaining observations in DLCC are depth-based kkNN and RF. In both cases, the temporary-cluster labels serve as the training labels, and we intentionally avoid additional hyperparameter tuning to keep the procedure lightweight and reproducible.
The kkNN option is a simple nonparametric classifier that can accommodate non-convex decision boundaries. Since DLCC already provides the depth-based similarity matrix $\mathbb{S}$, neighbors are obtained by sorting $\mathbb{S}_{i,\cdot}$, avoiding additional distance evaluations. Moreover, the choice of $k$ can be made data-adaptive via the notion of natural neighbors [47]. Concretely, we take the smallest $k$ such that at least $95\%$ of observations have at least one mutual neighbor within their top-$k$ neighbor lists, where the $95\%$ threshold serves as a safeguard against potential outliers.
As for RF, it is a widely used nonparametric classifier that makes no explicit shape assumptions and often remains effective in high-dimensional spaces [48]. Again, we do not perform hyperparameter tuning and fix the number of trees at $100$.
III-C Complexity Analyses
The overall computational cost of DLCC is dominated by constructing the depth-based similarity matrix $\mathbb{S}$ and the associated local-center estimation. As discussed in Section II, computing $\mathbb{S}$ in its exact form requires $O(n^{3})$ operations. Estimating local centers over $n$ neighborhoods of size $s$ incurs an additional cost of $O(ns^{2})$. Subsequent filtering and grouping steps operate only on the $T$ local-center candidates, where $T\ll n$, and thus contribute negligibly to the overall complexity.
The temporary clustering step and the depth-based kkNN classification are performed on top of the precomputed similarity matrix $\mathbb{S}$ and incur at most linear cost in $n$. For the RF option, training a forest with $100$ trees has a typical cost roughly $O(100\,n\log n)$.
Consequently, the overall complexity is dominated by the $O(n^{3})$ cost of constructing $\mathbb{S}$. When the data dimension $d$ is large, this complexity is expressed as $O(n^{3}d)$, since the computation of spatial depth involves the evaluation of $L^{2}$ norms of averaged unit vectors.
IV Exploring Parameter Selection
IV-A Strategy Chosen
Under the min strategy, DLCC attempts to find centroids of clusters from filtered centers. The variation in the number of filtered centers across different clusters allows for better flexibility than center-based clustering methods. The min strategy is more suitable when the sizes of all clusters are comparable. Moreover, because it groups centers more conservatively, it is often more reliable for overlapping clusters.
The max strategy views each cluster as a union of multiple finite convex regions, where each region is approximated by the convex hull of neighborhood points associated with a filtered center. Filtered centers are grouped via connectivity. Specifically, two filtered centers $c_{y}$ and $c_{z}$ are assigned to the same group if there exists a chain $c_{p_{1}},c_{p_{2}},\ldots,c_{p_{h}}$ with $c_{p_{1}}=c_{y}$, $c_{p_{h}}=c_{z}$, such that the similarity between neighborhoods of any adjacent pair in the chain exceeds $\delta$. Consequently, the max strategy is typically more effective for non-convex structures or highly unbalanced cluster sizes.
When no prior information about the dataset is available, we recommend selecting the strategy via visualization. For high-dimensional data, a 2D embedding (e.g., t-SNE [45]) is often informative: if clusters appear well separated and roughly balanced, the min strategy is recommended; otherwise, the max strategy is preferred.
IV-B Max Strategy